OpenStudy (anonymous):

A 1000-kg car traveling at 50 m/s runs into the rear of a stopped car that has a mass of 1500 kg and they stick together. What is the speed of the cars after the collision?

OpenStudy (asnaseer):

It would be better for you to post this in a more appropriate group (like Physics) and then post a link to it in any of the chat windows if you find no one available to help out in that group. This helps keep the groups on topic. Please refer to our Code of Conduct if you require further clarification: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Try the conservation of momentum. The sum of the momenta before the collision will be the momentum of the system after the collision. Momentum for an object of mass \(m\) and velocity \(\vec v\) is given by \(\vec p = m\vec v\).

OpenStudy (anonymous):

so this is a momentum problem! gotta love physics =) ok so usually we use something like this. p1=m1v1 p1=1000kg*50m/s p1=50000kg*m/s p2(p2=p1 because they are both one mass at the end of the question.) sooo 50,000=2500kg*v 50,000/2500=20m/s=v There ya go. =)